OpenStudy (gabbyalicorn):

What is the surface area of a rectangular prism with a length of 7 cm, a width of 10 cm, and a height of 2 cm?

OpenStudy (gabbyalicorn):

@amberosales @iGreen

OpenStudy (rainbow_rocks03):

What do you think the answer is @gabbyalicorn ?

OpenStudy (black_ninja123):

formula is 2(wl+hl+hw)

OpenStudy (gabbyalicorn):

2 x 70 + 14 + 20?

OpenStudy (rainbow_rocks03):

The area for the top and bottom side is l× w + l × w = 2 × l × w

OpenStudy (gabbyalicorn):

i thought black ninja explained tht

OpenStudy (gabbyalicorn):

i agree with u on tht 1

OpenStudy (rainbow_rocks03):

solve the one I just said .The area for the top and bottom side is l× w + l × w = 2 × l × w

OpenStudy (gabbyalicorn):

i dunno i tried

OpenStudy (gabbyalicorn):

i dont get it i did it wrong i gues

OpenStudy (rainbow_rocks03):

show me how you did it.

OpenStudy (gabbyalicorn):

2 x 70 + 14 + 20 = 174

OpenStudy (gabbyalicorn):

@rainbow_rocks03

OpenStudy (black_ninja123):

2(70+14+20) Now see what you get

OpenStudy (gabbyalicorn):

208

OpenStudy (black_ninja123):

Correct

OpenStudy (gabbyalicorn):

:D yay!